OpenStudy (shal):

anyone would like to help me with algebra 2?

OpenStudy (mathmale):

Please type in your problem first thing. Then, if you wish, ask, "anyone would like to help me with algebra 2?"

OpenStudy (anonymous):

What is your question?

OpenStudy (shal):

@Electric_Blue

OpenStudy (anonymous):

sorry i cant help you :'( never learned that or forgot it

OpenStudy (shal):

okay

OpenStudy (mathmale):

I assume that there are FOUR possible answer graphs, but you're sharing only one (now three). Before helping you, I'd ask that you first try graphing simpler functions such as these:\[f(a)=\frac{ 1 }{ a^2 }, g(a)=\frac{ 1 }{ a^4 }\] ... to get an idea of how these functions behave.

OpenStudy (mathmale):

Notice that both of these simpler functions are "even" functions, as is the given function. What implications does that have when you're graphing these functions?

OpenStudy (shal):

theres 4 now

OpenStudy (shal):

right?

OpenStudy (mathmale):

Yes. Notice that both of these simpler functions are "even" functions, as is the given function. What implications does that have when you're graphing these functions?

OpenStudy (shal):

all at 0.5?

OpenStudy (mathmale):

sorry, but I see no relevance of "all at 0.5?" to the problem at hand. can you graph f(a)=a^2?

OpenStudy (shal):

no

OpenStudy (mathmale):

Have you any idea at all what the shape of the graph would be? This is a very common, basic function, one you could ideally graph.

OpenStudy (mathmale):



OpenStudy (shal):

vice versa?

OpenStudy (mathmale):

Do you have a graphing calculator, and, if so, do you know how to graph functions with it?

OpenStudy (shal):

im not getting helped

OpenStudy (mathmale):

I am sorry you're not happy with this conversation, but at the same time I don't consider it my role to start you out from ground zero.

OpenStudy (shal):

no

OpenStudy (mathmale):

I'm not trying to make you feel bad, you understand, but I must tell you that you seem to lack the basic understanding of graphing necessary to do this problem you've posted here. Are you home schooled? or taking an algebra course online? or what?

OpenStudy (mathmale):

Look: I have given you the graph of y=x^2. The graph of the reciprocal of this function, \[f(x)=\frac{ 1 }{ x^2 }\] looks liike the following: